OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x – 6y = –15 x – 3y = –3 Answer (–6y – 15) – 3y = –3 x – 3(6y – 15) = –3 (6y – 15) – 3y = –3 x – 3(–6y – 15) = –3

OpenStudy (anonymous):

(6y – 15) – 3y = –3

OpenStudy (anonymous):

x – 6y = –15 rearranging the first equation, we get x = -15 + 6y or we can say x = 6y - 15 substituting x = 16y - 15 in second equation u get (6y - 15) – 3y = –3